A 0.73 kg magnetic target is suspended on a string. A 0.025 kg magnetic dart, shot
horizontally, strikes it head on. The dart and the target together swing up 12 cm above the
initial level. What was the initial velocity of the dart?
1
Expert's answer
2011-08-25T13:28:04-0400
Let m1 be the mass of the magnetic dart, m2 - mass of the target.
V - initial velocity of the dart.
Consider the target and the dart didn't have the initial potential energy,
let's use the law of conservation of energy:


E1 = E2,
where
E1 - initial energy of the system,
E2 - energy of the system after the collision.
E1 = Ek - kinetic energy,
E2 = Ep - potential energy


m1*V2 = (m1 + m2)*g*h


V = sqrt((m1+m2)*g*h/m1)
V = sqrt((0.025+0.73)*9.8*0.12/0.025) = 5.96 m/s.


So, the initial velocity of the dart is 5.96 m/s.